# ABScribe: Rapid Exploration of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models

### v1

@@ Abstract.
Exploring alternative ideas by rewriting text is integral to the writing process. State-of-the-art large language models (LLMs) can simplify writing variation generation. However, current interfaces pose challenges for simultaneous consideration of multiple variations: creating new versions without overwriting text can be difficult, and pasting them sequentially can clutter documents, increasing workload and disrupting writers’ flow. To tackle this, we present ABScribe, an interface that supports rapid, yet visually structured, exploration of writing variations in human-AI co-writing tasks. With ABScribe, users can swiftly produce multiple variations using LLM prompts, which are auto-converted into reusable buttons. Variations are stored adjacently within text segments for rapid in-place comparisons using mouse-over interactions on a context toolbar. Our user study with 12 writers shows that ABScribe significantly reduces task workload (), enhances user perceptions of the revision process () compared to a popular baseline workflow, and provides insights into how writers explore variations using LLMs.
@@ 1. Introduction
Revision is an essential part of the writing process (Zinsser, 2006; Hemingway, 2014; Goldberg, 2016; McPhee, 2017). Professional writers often write and rewrite text hundreds of times (Sommers, 1980; Chenoweth, 1987) and recommend rewriting as a core strategy for writing well (Zinsser, 2006; Chenoweth, 1987). Effective revision goes beyond minor editorial changes, and may help writers rework ideas, and powerfully affect their knowledge (Fitzgerald, 1987; Sommers, 1980) as they explore alternative variations to find a line of argument (Fitzgerald, 1987). The revision process is iterative (Du et al., 2022; Fitzgerald, 1987): happens in repeated cycles, throughout the writing process, granular (Seow, 2002; McPhee, 2017): happens at the word, sentence, or paragraph-level, and non-linear(Cheung, 2016; Sommers, 1980): requires constant reconsideration of potential variations of existing text elements throughout the passage. Current writing interfaces tend not to support a non-linear revision process and predominantly support linear representations of revision history (e.g. revision history in popular word processing software such as Google docs or Microsoft Word). While these tools do support iterative and granular edits, it remains difficult for writers to simultaneously consider multiple writing variations and to organize them without replacing earlier text, or cluttering documents when writers resort to pasting them in sequence. Insights from HCI and traditional design practice suggest that simultaneous consideration of multiple (at least 5) variations can lead to better ideas (Tohidi et al., 2006) and avoid fixation (Jansson and Smith, 1991; Tohidi et al., 2006; Greenberg and Buxton, 2008; Buxton, 2010). We hypothesize that this may apply to writing, provided that writers are given adequate support in managing multiple variations with minimal workload.
@@
Advanced Large Language Models (LLMs) such as ChatGPT (cha, 2022), GPT-4 (OpenAI, 2023), PaLM 2 (PaL, 2023), and LLaMA 2 (Touvron et al., [n. d.]), can enable writers to generate multiple variations of text via prompting (Brown et al., 2020; Dale, 2021), potentially reducing the workload of generating text variations. However, easier generation can exacerbate challenges surrounding the systematic storage, comparison, and modification of multiple variations with existing chat-based and in-place editing interfaces where users are required to find text variations in linear chat histories or store them in their document editor as comments or separate in-line text blocks. As text variations become easier to generate using AI, they become harder to manage.
@@
Recent HCI studies on LLM-based tool design have mainly focused on prompt engineering (Brade et al., 2023; Zamfirescu-Pereira et al., 2023; Giray, 2023) and exploring the generative capabilities of LLMs (Yuan et al., 2022; Lee et al., 2022b; Wu et al., 2023; MacNeil et al., 2023). For example, Zamfirescu-Pereira et al. investigated ways to support non-AI experts with crafting effective prompts, and Yuan et al. (Yuan et al., 2022) explored how users might use LLMs for creative writing. They found that the output of the model did not need to be perfect to be useful to users. Many users found the output useful, even if they had to significantly revise the text or chose not to incorporate it into their final draft (Yuan et al., 2022). This highlights the potential value of designing affordances that help manage imperfect AI-generated variations. Even if these variations don’t make it into the final text, they might still be valuable to consider.
@@
In this paper, we present ABScribe111We name our system ABScribe to reference how we label multiple variations using the alphabet. Note that we support variations beyond just A and B. The Hover buttons can include multiple variations: A, B, C, D, E, etc.–a novel writing interface that supports the rapid exploration of multiple writing variations in LLM-based human-AI co-writing tasks. We draw inspiration from Kim et al. design framework which shows the potential for object-oriented interactions with LLMs to encourage iteration and experimentation during writing (Kim et al., 2023), and propose a suite of five interface elements that support writers in swiftly exploring multiple writing variations by interacting with an ensemble of five interface elements: (i) Variation Components that store multiple human and AI-generated variations within flexible text segments in a non-linear manner, without overwriting text; (ii) Hover Buttons that reveal corresponding versions inside a Variation Component when users hover their mouse over them, allowing for rapid comparisons without breaking text flow; (iii) the Variation Accordion that organizes all variations in a navigable format; (iv) AI Buttons that automatically encapsulates LLM instructions into reusable buttons that can be applied across different Variation Components; and (v) AI Insert that allows writers to insert LLM-generated text directly into the document (see Figure 1).
@@
To validate our design, we conducted a controlled evaluation study and interviews with 12 writers comparing ABScribe with a widely-used baseline worklow consisting of an AI integrated rich text editor based on GPT-4, with a chat-based AI assistant. Our findings demonstrate that ABScribe significantly reduces subjective task workload (), and enhances user perceptions of the revision process (), compared to the baseline. The key contributions of our work are as follows:
@@
The design and implementation of ABScribe, an LLM-enhanced writing interface that supports the rapid exploration of multiple text variations in human-AI co-writing tasks.
@@ 2. Related Work
We review literature in HCI and traditional design practices on weighing multiple alternatives, and discuss the relevance of this design method when revising writing, guided by they theory on revision process in writing. We delve into difficulties that arise when trying to support multiple variation exploration using existing editing interfaces and contrast between chat-based and in-place interfaces to situate our design within a broader class of Human-AI writing interfaces.
@@ 2.1. Exploring Multiple Variations
HCI and traditional design practice encourages the parallel exploration of multiple variations to help avoid fixation on a singular idea (Jansson and Smith, 1991; Goldschmidt, 2011), to reduce the chances of eliminating rough but innovative ideas due to premature evaluation (Greenberg and Buxton, 2008; Buxton, 2010), and to make us less prone to inflated subjective appraisals by giving us an opportunity to critically assessing ideas in relation to each other (Tohidi et al., 2006; Buxton, 2010). In this paper, we hypothesize that such parallel exploration of multiple variations may apply to the revision process during writing. Much like how a naive, linear implementation of an iterative design approach encourages the sequential refinement of ideas, when writers don’t have a way to organize and work with multiple text variations, they may end up committing to ideas too early, and focusing too much on surface level edits to refine their draft.
This is problematic because when we turn to research on the revision and the writing process, we see that revision goes beyond surface level edits (Fitzgerald, 1987), encompassing deeper writing subprocesses such as revising and evaluating ideas (Flower and Hayes, 1981a) and meaning discovery (Sommers, 1980). Experienced writers treat revision as a recursive, non-linear process (Sommers, 1980; Seow, 2002), and engage with the text in repeated cycles, with multiple objectives including finding the form or shape of an argument (Sommers, 1980), experimenting with vocabulary and style (Huff, 1983), and going back and forth between multiple composing activities as writers revise text (Faigley and Witte, 1981).
In addition to the rich-body of work underscoring the important and complex role of revision in writing, researchers have explored the benefits of adopting design language in writing pedagogy, such as characterizing writing pedagogy as a wicked (Rittel, 1967) design thinking problem (Leverenz, 2014; Purdy, 2014). There has also been some valuable work in HCI to support novel editing practices, such as supporting constraints and consistency in maintaining domain-specific terms across complex documents (Han et al., 2020a) using persistent, reified (Beaudouin-Lafon and Mackay, 2000) text selections, and present the idea and implementation of a varientlet, that allows writers to store and compare two variations. However, further innovation in this space is needed to design affordances for writers that support the simultaneous consideration of multiple variations during the revision process.
In this paper, we contribute to this line of work, and present a suite of interface elements that work together in supporting writers with the rapid exploration of multiple text variations in a non-linear fashion, in alignment with the nature of the revision process, and offer empirical insights into the applicability of design ideas in HCI on the parallel consideration of variations to the specific task of revision in writing.
@@ 2.2. Working with Multiple Variations from Large Language Models
As we work toward leveraging advanced LLMs such as ChatGPT (cha, 2022), GPT-4 (OpenAI, 2023), PaLM 2 (PaL, 2023), and LLaMA 2 (Touvron et al., [n. d.]), which can enable writers to generate multiple variations of text based on different parts of their writing using prompts, HCI researchers looking to design writing interfaces are faced with several challenges. These include dealing with the non-deterministic nature of these models (Kim et al., 2023; Jiang et al., 2022), systematically exploring their capabilities (Yuan et al., 2022; Lee et al., 2022b), and making prompt-writing easier for AI-novices (Zamfirescu-Pereira et al., 2023; Shakeri et al., 2021).
There has also been work on managing the output from generative AI to support the exploration of variations in different contexts such as exploring images (Brade et al., 2023; Lee et al., 2022a), and multi-modal interactions beyond text prompts to explore generative AI (Liu, 2023). However, our understanding of how to best organize the prolific output in AI-augmented writing workflows is still limited.
We draw inspiration from Kim et al’s work on the use of object-oriented interactions and reification (Han et al., 2020b) to encourage writers’ experimentation of LLM output, as well as prior work on revision control in writing (Han et al., 2020a; Han, 2020) and other domains (Head et al., 2019), and offer the design and implementation of a novel interface that tackles the problem of how to effectively organize multiple variations from an LLM in a way that minimizes task workload while supporting parallel exploration.
@@ 2.3. Chat-Based and In-Place Human-AI Co-Writing Interfaces
To help ground our interface design, we distinguish between two types of Human-AI Co-Writing interfaces into two types: conversational interfaces such as ChatGPT and Bard, and In-Place interfaces that directly inserts or modifies text in the document.
Chat-Based Interfaces: Currently the dominant mode, chat-based interfaces, like ChatGPT (cha, 2022), Bing Chat (bin, 2013), and Bard (Pichai, 2023), have gained immense popularity. These conversational interfaces are highly intuitive, and mimic human-to-human chat interactions, but lack scaffolding for crafting prompts, which can be difficult for novice AI users (Zamfirescu-Pereira et al., 2023; Kroll and Reid, 1994). Another significant limitation is the linear chat-log structure. In contrast to the non-linear nature of how revision happens in writing (Sommers, 1980; Fitzgerald, 1987) the text-variations generated using a chat-based interface are buried within linear chat-logs, impeding parallel exploration of multiple variations in-place, where the writer is editing the document.
In-Place Editing Interfaces: This type offers closer integration between the human and AI writer during the text editing process by adopting a more What You See Is What You Get (WYSIWYG) (Bly, 2007) approach where AI-generated text modifies the human text and vice versa. This offers increased flexibility over the edited content compared to chat and form-based interfaces by allowing users to edit individual sections of the text.
Recent research prototypes for LLM writing tools such as Wordcraft (Yuan et al., 2022) and CoAuthor (Lee et al., 2022b) allow for in-place editing. Wordcraft (Yuan et al., 2022) melds an in-text interface with diffrent options for users to continue a narrative based on prior text and replacing text selections with AI-modified content. CoAuthor explores GPT-3’s capabilities by capturing deep interactions between writers and GPT-3 via a similar in-place editing interface where users receive multiple edit suggestions from the model. Commercial tools like Grammarly222grammarly.com and Wordtune333wordtune.com also allow users to enhance their writing by revising text using AI-driven suggestions. However, once generated by the AI and modified by the user, previous text can become obscured or lost as older text is superseded by new edits. Even if text is auto-saved, it is often preserved as linear version histories, as is the case in Google Docs (an online word processor), or undo/redo histories, making it difficult to work with multiple text variations concurrently.
In our design, we adopt an in-place editing interface in a GPT-4 powered research prototype, offering a solution to overcome challenges surrounding the management of multiple text variations in human-AI co-writing tasks. We carefully construct a baseline interface that represents current workflows, providing fresh empirical insights based on our interviews with writers. These insights help us understand user perceptions of the revision process and explore how differences between in-place editing and chat-based AI writing companions impact their workflow.

### v3

# ABScribe: Rapid Exploration & Organization of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models

@@ Abstract.
Exploring alternative ideas by rewriting text is integral to the writing process. State-of-the-art Large Language Models (LLMs) can simplify writing variation generation. However, current interfaces pose challenges for simultaneous consideration of multiple variations: creating new variations without overwriting text can be difficult, and pasting them sequentially can clutter documents, increasing workload and disrupting writers’ flow. To tackle this, we present ABScribe, an interface that supports rapid, yet visually structured, exploration and organization of writing variations in human-AI co-writing tasks. With ABScribe, users can swiftly modify variations using LLM prompts, which are auto-converted into reusable buttons. Variations are stored adjacently within text fields for rapid in-place comparisons using mouse-over interactions on a popup toolbar. Our user study with 12 writers shows that ABScribe significantly reduces task workload (), enhances user perceptions of the revision process () compared to a popular baseline workflow, and provides insights into how writers explore variations using LLMs.
@@ 1. Introduction
Revision is essential to the writing process (Zinsser, 2001; Hemingway, 2014; Goldberg, 2016; McPhee, 2017). Professional writers often write and rewrite text hundreds of times (Sommers, 1980; Chenoweth, 1987) and recommend rewriting as a core strategy for writing well (Zinsser, 2001; Chenoweth, 1987). Effective revision goes beyond minor editorial changes, and may help writers rework ideas, and powerfully affect their knowledge (Fitzgerald, 1987; Sommers, 1980) as they explore alternative variations to find a line of argument (Fitzgerald, 1987). The revision process is iterative (Du et al., 2022; Fitzgerald, 1987): happens in repeated cycles, throughout the writing process, granular (Seow, 2002; McPhee, 2017): happens at the word, sentence, or paragraph-level, and non-linear(Cheung, 2016; Sommers, 1980): requires constant reconsideration of potential variations of existing text throughout the passage.
Current writing interfaces tend not to support a non-linear revision process and predominantly support linear representations of revision history (e.g., revision history in popular word processing software such as Google docs or Microsoft Word). While these tools do support iterative and granular edits, it remains difficult for writers to simultaneously consider multiple writing variations and to organize them without replacing earlier text, or cluttering documents when writers resort to pasting them in sequence. Insights from HCI and traditional design practice suggest that simultaneous consideration of multiple (at least 5) variations can lead to better ideas (Tohidi et al., 2006) and avoid fixation (Jansson and Smith, 1991; Tohidi et al., 2006; Greenberg and Buxton, 2008; Buxton, 2010). We hypothesize that this may apply to writing, provided that writers are given adequate support in managing multiple variations with minimal workload.
@@
Advanced Large Language Models (LLMs) such as ChatGPT (OpenAI, 2022), GPT-4 (OpenAI, 2023), PaLM 2 (Google, 2023), LLaMA 2 (Touvron et al., 2023), and Gemini(Pichai and Hassabis, 2023) can enable writers to generate multiple variations of text via natural language prompting (Brown et al., 2020; Dale, 2021; Gero et al., 2022; Min et al., 2023), potentially reducing the workload of generating text variations. However, easier generation can exacerbate challenges surrounding the systematic organization, comparison, and modification of multiple variations with existing chat-based and in-place editing interfaces where users are required to find text variations in linear chat histories and store them in their document as comments or separate in-line text blocks. As text variations become easier to generate using AI, they become harder to manage.
@@
Recent HCI studies on LLM-based tool design have mainly focused on prompt engineering (Brade et al., 2023; Zamfirescu-Pereira et al., 2023; Liu and Chilton, 2022; Jiang et al., 2022; Wu et al., 2022; Dang et al., 2023) and exploring the generative capabilities of LLMs (Yuan et al., 2022; Lee et al., 2022a; MacNeil et al., 2023; Meyer et al., 2022; Mirowski et al., 2023). For example, Zamfirescu-Pereira et al. investigated ways to support non-AI experts with crafting effective prompts, and Dang et al. (Dang et al., 2023) found that users preferred choosing from multiple suggestions and using content from the draft to guide LLM usage rather than writing external prompts which require more effort. Yuan et al, (Yuan et al., 2022) explored how users might use LLMs for creative writing and found that the output of the model did not need to be perfect to be useful to users. Many users found the output useful, even if they had to significantly revise the text or chose not to incorporate it into their final draft (Yuan et al., 2022). This highlights the potential value of designing affordances that help manage imperfect AI-generated variations. Even if these variations don’t make it into the final text, they might still be valuable to consider.
@@
In this paper, we present ABScribe111We name our system ABScribe to reference how we label variations using the alphabet. Note that the interface supports variations beyond just A and B. The Popup Toolbar can include multiple variations: A, B, C, D, E, etc.–a novel writing interface that supports the rapid exploration and organization of multiple writing variations in LLM-based human-AI co-writing tasks. We draw inspiration from a design framework by Kim et al. which shows the potential for object-oriented interactions with LLMs to encourage iteration and experimentation during writing (Kim et al., 2023a), and propose an ensemble of five interface elements that support writers in swiftly working with multiple writing variations: (i) Variation Fields that store multiple human and AI-generated variations within flexible text segments in a non-linear manner, without overwriting text; the (ii) Popup Toolbar that reveal corresponding variations inside a Variation Field when users hover their mouse over buttons representing each variation, allowing for rapid comparisons without breaking text flow; (iii) the Variation Sidebar that organizes all variations in a navigable format within an accordion UI; (iv) AI Modifiers that automatically encapsulates LLM instructions into reusable buttons that can be applied across different Variation Fields; and (v) the AI Drafter which allows writers to insert LLM-generated text directly into the document (see Figure 1).
@@
To validate our design, we conducted a controlled evaluation study involving revision tasks and interviews with 12 writers comparing ABScribe with a widely-used baseline worklow consisting of an AI integrated rich text editor based on GPT-4, with a chat-based conversational AI assistant. Our findings demonstrate that ABScribe significantly reduces subjective task workload (), and enhances user perceptions of the revision process (), compared to the baseline. The key contributions of our work are as follows:
@@
The design and implementation of ABScribe, an LLM-enhanced writing interface that supports the rapid exploration and organization of multiple text variations in human-AI co-writing tasks. We have made the full code for ABScribe open source and accessible via GitHub222Link to GiHub Repository: https://github.com/mohireza/abscribe/.
@@ 2. Related Work
We first review literature in HCI and traditional design practices on considering multiple alternatives, and discuss how this design approach is pertinent when revising text, informed by the theory on the revision process in writing (section 2.1). Then, we examine this work’s alignment with the rapidly growing research in AI-assisted writing, focusing on the effects of exposing writers to diverse AI-generated suggestions, and explore some challenges that arise when trying to support multiple variation editing using existing interfaces (section 2.2). Finally, we contrast between chat-based and in-place interfaces to situate our design within a broader class of Human-AI writing interfaces (section 2.3).
@@ 2.1. Exploring Multiple Variations
HCI and traditional design practices encourage the parallel exploration of multiple variations to help avoid fixation on a singular idea (Jansson and Smith, 1991; Goldschmidt, 2011), to reduce the chances of eliminating rough but innovative ideas due to premature evaluation (Greenberg and Buxton, 2008; Buxton, 2010), and to make us less prone to inflated subjective appraisals by giving us an opportunity to critically assess ideas in relation to each other (Tohidi et al., 2006; Buxton, 2010). In this paper, we hypothesize that such parallel exploration of multiple variations may apply to the revision process during writing. Much like how a naive, linear implementation of an iterative design approach encourages the sequential refinement of ideas, when writers do not have a way to organize and work with multiple text variations, they may end up committing to ideas too early, and focusing too much on surface level edits to refine their draft.
This is problematic because when we turn to research on the revision and writing process, we see that revision goes beyond surface level edits (Fitzgerald, 1987), encompassing deeper writing subprocesses such as revising and evaluating ideas (Flower and Hayes, 1981a) and meaning discovery (Sommers, 1980). Experienced writers treat revision as a recursive, non-linear process (Sommers, 1980; Seow, 2002), and engage with the text in repeated cycles, with multiple objectives including finding the form or shape of an argument (Sommers, 1980), experimenting with vocabulary and style (Huff, 1983), and going back and forth between multiple composing activities during revision (Faigley and Witte, 1981).
In addition to the rich-body of work underscoring the important and complex role of revision in writing, researchers have explored the benefits of adopting design language in writing pedagogy, such as characterizing writing pedagogy as a wicked (Rittel, 1967) design thinking problem (Leverenz, 2014; Purdy, 2014). There has also been some valuable work in HCI in designing novel editing practices, such as supporting constraints and consistency in maintaining domain-specific terms across complex documents (Han et al., 2020a) using persistent, reified (Beaudouin-Lafon and Mackay, 2000) text selections that allow writers to store and compare two variations. However, further innovation in this space is needed to design affordances for writers that support the simultaneous consideration of multiple variations during the revision process.
We contribute to this line of work, and present a suite of interface elements that work together in supporting writers with the rapid exploration and organization of multiple text variations in a non-linear fashion, in alignment with the nature of the revision process, and offer empirical insights into the applicability of design ideas in HCI on the parallel consideration of variations to the specific task of revision in writing.
@@ 2.2. Working with Multiple Variations from Large Language Models
There has been an explosion of recent work (Gero et al., 2022; Giray, 2023; Dang et al., 2023; Mirowski et al., 2023; Buschek et al., 2021; Gero et al., 2023; Long et al., 2023; Chung et al., 2022; Singh et al., 2023; Kim et al., 2023b, a; Lee et al., 2022a; Yuan et al., 2022; Biermann et al., 2022; Shakeri et al., 2021; Goodman et al., 2022) on interfaces for writing with LLMs in tandem with rapid advancements in AI writing capabilities using LLMs (Min et al., 2023; Vaswani et al., 2017; Radford et al., 2018; Devlin et al., 2018; Brown et al., 2020), examining various aspects of the AI-assisted writing process including prompting (Wu et al., 2022; Giray, 2023), text suggestions (Dang et al., 2023; Buschek et al., 2021), creative ideation (Gero et al., 2022; Chung et al., 2022), and social dynamics (Biermann et al., 2022; Gero et al., 2023). Among these, prior studies on the impact of showing multiple text suggestions from the AI (Buschek et al., 2021; Dang et al., 2023) are of particular relevance to our design. Buschek et al. (Buschek et al., 2021) highlighted the benefits of multiple parallel suggestions for ideation but warned against the simultaneous cost to efficiency. Dang et al. (Dang et al., 2023) coined and distinguished between diegetic prompts that are part of draft with non-diegetic prompts that are external instructions to the LLM. They uncovered user behavioural patterns indicating a tendency to choose from multiple suggestions rather than writing non-diegetic prompts as that requires more effort, and found that users prefer to use the draft, i.e., diegetic informtion, to guide LLM prompts.
AI Modifiers, for instance, simplify prompt-writing by turning modification instructions into reusable buttons, working closely with Variation Fields to offer users multiple options and utilize diegetic information directly from the draft. We also incorporate Kim et al.’s concepts of object-oriented interactions with LLMs(Han et al., 2020b) and draw from existing revision control research (Han et al., 2020a; Han, 2020; Head et al., 2019) to create interface elements that effectively manage multiple LLM variations, reducing task workload and facilitating parallel exploration.
@@ 2.3. Chat-Based and In-Place Human-AI Co-Writing Interfaces
To ground our interface design, we distinguish between two types of Human-AI Co-Writing interfaces: Chat-Based interfaces such as ChatGPT and In-Place interfaces that directly insert or modify text in a document.
Chat-Based Interfaces: Currently the dominant mode, chat-based interfaces, like ChatGPT (OpenAI, 2022), Bing Chat (Microsoft, 2013), and Bard (Pichai, 2023), have gained immense popularity. These interfaces are highly intuitive, and mimic conversational interactions between humans, but lack scaffolding for crafting prompts, which can be difficult for novice AI users (Zamfirescu-Pereira et al., 2023; Kroll and Reid, 1994). In the context of writing with AI, simple turn-taking between the writer and the AI does not provide adequate support to writers in steering the iterative output from the LLM (Chung et al., 2022).
Another significant limitation is the linear chat-log structure. In contrast to the non-linear nature of how revision happens in writing (Sommers, 1980; Fitzgerald, 1987), the text-variations generated using a chat-based interface are buried within linear chat-logs, impeding parallel exploration of multiple variations in-place, where the writer is editing the document.
In-Place Editing Interfaces: This type offers closer integration between the human and AI writer during the text editing process by adopting a more What You See Is What You Get (WYSIWYG) (Bly, 2007) approach where AI-generated text modifies the human text and vice versa. This offers increased flexibility over the edited content compared to chat-based interfaces by allowing users to edit individual sections of the text.
Recent research prototypes like Wordcraft (Yuan et al., 2022) and CoAuthor (Lee et al., 2022a), as well as commercial tools such as Grammarly333grammarly.com, Notion AI, and Wordtune444wordtune.com, offer in-place editing features for LLM writing assistance. Wordcraft (Yuan et al., 2022) combines an editing interface with options for narrative continuation and AI-modified text replacement, while CoAuthor (Lee et al., 2022a) leverages GPT-3 for deep interaction and multiple edit suggestions. However, these tools typically overwrite previous text with new edits, obscuring the original content. Saved versions are often linear, as seen in Google Docs, or limited to undo/redo histories, complicating the management of multiple text variations simultaneously.
In our design, we adopt an in-place editing interface in a GPT-4 powered research prototype, offering a solution to overcome challenges surrounding the management of multiple text variations in human-AI co-writing tasks. We provide fresh empirical insights from interviews with writers on how our in-place interface design compares with existing chat-based writing workflows, and their impact on user perceptions of the revision process and task workload.